differentiate the following function with respect to x
(log x)^x

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
y=logx ^x
log(y)= x log(log(x))
now taking derivative
'1/y*dy/dx = log(logx)+x/xlog(x)= log(logx)-logx
dy/dx=y (log(logx)-logx)